OpenStudy (anonymous):

Math Analysis: Solve for 0

OpenStudy (anonymous):

3pie/4 and 7pie/4

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

try it this way... sinx + cosx = 0 sinx = -cosx sinx/cosx = -1 tanx = -1 now look at the unit circle (or table of trig values or calculator) that solves this last equation.

OpenStudy (anonymous):



OpenStudy (anonymous):

ty :3

OpenStudy (anonymous):

@obada it was in radians not in pie btw =P

OpenStudy (anonymous):

ok ., in x= 135/180 pie , x= 315/180pie

OpenStudy (anonymous):

other method is that in the second quadrant and 4th quadrant, you have sinx and cosx in oppsite signs. Since sinx and cosx have equal magnitudes only at pie/4 so in second quadrant x= 3pie/4 and in 4th quadrant x=7pie/4